Session: Feb 1998 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (tony foley at Tue Feb 24 14:56:38 MET 1998) Minor failures
First problems; jump in formatter clock at about 1128UT (GPS-fmout changes from 2.1 to 3.1 microseconds). The telescope control computer was rebooted at about 12:17UT, so that when it restarted at 12:23UT the telescope was ~0.06 degrees offsource (telescopes try to follow using their microprocessors).

Torun (K. Borkowski at Tue Feb 24 16:17:16 MET 1998) Success
No known problems

Onsala (A. Polatidis at Thu Feb 26 13:03:33 MET 1998) Success
The inchworm motor for HEAD2 (the READ head) is not working. HEAD 1 (the WRITE head) is working properly so the data shound be OK. Please ignore the results of the readback tests in the log.

Jodrell (Alastair Gunn at Thu Feb 26 14:59:16 MET 1998) Minor failures
Strong RFI - receiver often saturated. On source 3C345 at 09:02 UT. No other known problems.

Calibration (jean\-francois Desmurs at Mon Mar 2 18:52:29 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name gc20b.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Thu Mar 5 16:12:14 MET 1998) Success